For dinner, Marcus consumes 175 grams of carbohydrate, 60 grams of protein, and 20 grams of fat. In addition, Marcus decides that he wants a glass of wine with his meal. If he drinks one glass of wine containing 8 grams of alcohol, how many total kilocalories (kcal) does he consume in this meal?
Answer #1

Calorie from carbo hydrates= 175× 4= 700 Cal

Calorie from protein =60×4 = 240 Cal

Calorie from fat = 20×9 = 180 Cal

A usual glass (4 ounce) of wine provides 100 Calories

Alcohol content do not let out any calories.

Total Calories in his meal = 700+ 240+ 180+ 100 = 1220 Calories = 1.22 KCal